### v1

# Path Evolution Model for Endogenous Channel Digital Twin towards 6G Wireless Networks

Abstract
Massive Multiple Input Multiple Output (MIMO) is critical for boosting 6G wireless network capacity. Nevertheless, high dimensional Channel State Information (CSI) acquisition becomes the bottleneck of 6G massive MIMO system. Recently, Channel Digital Twin (CDT), which replicates physical entities in wireless channels, has been proposed, providing site-specific prior knowledge for CSI acquisition. However, external devices (e.g., cameras and GPS devices) cannot always be integrated into existing communication systems, nor are they universally available across all scenarios. Moreover, the trained CDT model cannot be directly applied in new environments, which lacks environmental generalizability. To this end, Path Evolution Model (PEM) is proposed as an alternative CDT to reflect physical path evolutions from consecutive channel measurements. Compared to existing CDTs, PEM demonstrates virtues of full endogeneity, self-sustainability and environmental generalizability. Firstly, PEM only requires existing channel measurements, which is free of other hardware devices and can be readily deployed. Secondly, self-sustaining maintenance of PEM can be achieved in dynamic channel by progressive updates. Thirdly, environmental generalizability can greatly reduce deployment costs in dynamic environments. To facilitate the implementation of PEM, an intelligent and light-weighted operation framework is firstly designed. Then, the environmental generalizability of PEM is rigorously analyzed. Next, efficient learning approaches are proposed to reduce the amount of training data practically. Extensive simulation results reveal that PEM can simultaneously achieve high-precision and low-overhead CSI acquisition, which can serve as a fundamental CDT for 6G wireless networks.
I Introduction
Massive Multiple Input Multiple Output (MIMO) is pivotal for 6G wireless networks, including ultra-high spectral efficiency, massive access and immersive sensing capabilities [1]. To fully leverage large-scale antenna arrays, Channel State Information (CSI) knowledge is critical for beamforming and precoding operations in massive MIMO systems. However, CSI acquisition overhead also rapidly increases in large antenna arrays [2]. Consequently, effective throughput will not consistently grow with the number of antennas.
High dimensional CSI acquisition becomes an inevitable challenge to meet promising 6G applications. To support numerous 6G applications with high throughput, there are two main requirements for CSI acquisition: (1) high CSI precision to facilitate beamforming and precoding operations; (2) low acquisition overhead for simultaneously serving numerous User Equipments (UEs).
I-A Channel Digital Twin
To resolve the bottleneck of high dimensional CSI acquisition, mainstream approaches directly utilize the measured CSI for prediction, which can also be enhanced by deep learning [3]. However, their channel acquisition accuracy dramatically relies on pilot densities, which cannot achieve high CSI precision when pilot densities are low. Additionally, deep learning related CSI extrapolation and interpolation heavily depend on the training dataset, whose generalization capability cannot be guaranteed in dynamic environments.
Recently, Channel Digital Twin (CDT) has emerged as an effective tool for CSI acquisition. CDT endeavors to reflect physics entities in wireless channels with digital representations, which offers site-specific prior knowledge for CSI acquisition. Compared to the aforementioned pure pilot-based approaches, CDT can greatly reduce pilot overhead to attain high-precision CSI, which offers an additional dimension for CSI acquisition. Currently, there are two main types of CDTs. One is vision information, which provides geometric and object information to attain the properties of wireless channel [4, 5]. Another is Channel Knowledge Map (CKM), which serves as a channel property database tagged with UE location [6]. To infer wireless channel property (e.g., beam index and path parameters) with input data (e.g., images and UE locations), deep learning is usually adopted in current CDTs.
Despite the prior knowledge, current CDTs still encounter two challenges. On the one hand, current CDTs cannot be instantly integrated into existing communication systems due to their nature of exogeneity. Once the extrinsic input outside of communication systems (e.g., images from cameras and locations from GPS devices) is unavailable, the operation of current CDTs will be interrupted. For instance, the operation can be challenging for indoor UEs due to occlusion and weak indoor GPS signals. On the other hand, learning of current CDTs is environment-dependent. Explicitly, neural networks in current CDTs need to be retrained when operating in new environments, which induces a huge data collection and model retraining burden for large-scale deployment.
In summary, CDT is an effective tool to enable superior applications for 6G wireless networks. However, when considering the aforementioned two challenges, operation of current CDTs is still challenging. Thus, the key question is can we find a new type of CDT that is endogenous within communication systems and is environment-generalizable?
I-B Proposed Path Evolution Model
To this end, Path Evolution Model (PEM) is proposed as a novel CDT to meet the goals of intra-system endogeneity and environmental generalizability. As shown in the top left of Fig. 1, the wireless channel between Base Station (BS) and UE is composited by multiple paths, which are basic physical entities for transmission, reflection and scattering. Due to UE mobility, feature (e.g., delay, Angle of Arrival (AoA), power) of each physical path consistently evolves with time, as shown in the top right of Fig. 1. Since channel measurements are frequently conducted in existing communication systems, as shown at the bottom of Fig. 1, the concept of PEM aims to extract each physical path and then attain digital replicas of their feature temporal evolutions from the consecutive channel measurements. Then, CDT is yielded when digital replicas of all path evolutions are summed up. Thus, the constructed PEM can be applied to directly attain the desired path feature at any instant, which greatly reduces pilot overhead and enables the functionality of CDT. Firstly, the processing of PEM only requires existing channel measurements from communication systems, which is fully endogenous. Secondly, path feature evolutions in PEM are determined by universal electromagnetic (EM) propagation principles and regular user mobility. As a result, a generalizable path feature evolution pattern can be assumed among dynamic environments. Therefore, when a neural network is trained to fit target path feature evolution, it can generalize to different environments as well.
Figure 1: Structure of PEM, where digital replicas of physical path evolutions are created.
I-C Advantages
Proposed PEM exhibits three inherent advantages over other existing CDTs:
  1. 1.
    Fully-endogenous: The intra-system endogeneity of PEM has two folds of virtues. On the one hand, only existing channel measurements within communication systems are required, which is free of other external devices. On the other hand, communication signals can realize larger coverage than either light or GPS signals. As a result, the outage probability of PEM can be much lower than the existing vision information-based and CKM-based CDTs.
  2. 2.
    Self-sustaining: PEM can effectively adapt to dynamic channels, where the digital path feature evolution can be equivalently viewed as a function of time and the explicit expression is controlled by historical path features. Since channel measurements are consecutive during the operation, the historical path features are progressively updated. In this way, the constructed path feature evolution is consistently refreshed from the latest channel measurements, which can minimize the error between digital and physical path evolution in real time.
  3. 3.
    Environment-generalizable: Trained neural network in PEM shows remarkable environmental generalizability since it endeavors to fit the environment-generalizable target path feature evolution. Such environmental generalizability greatly relieves the data collection and model retraining burden in new environments, which enables large-scale PEM deployment in 6G wireless networks.
I-D Our Contributions
In this paper, we propose the concept of PEM, aiming to create digital replicas of physical path evolutions upon communication systems. Firstly, we propose an intelligent operation framework to efficiently construct, maintain and apply PEM in dynamic wireless channels, which is compatible with existing communication systems and can be operated at low cost. Secondly, the environmental generalization capability of PEM is carefully analyzed and compared with existing CDT learning. Thirdly, efficient learning approaches are proposed to maximize the robustness of PEM under practical data collection constraints. In the numerical simulations, a good match can be found by comparing the physical and digital path evolutions, which validate the functionality of the proposed PEM. To justify the pilot overhead reduction capability of PEM, channel prediction under different time-domain pilot densities is investigated. Compared to CSI acquisitions without CDT, both high-precision and low-overhead CSI acquisition can be achieved with PEM. In addition, PEM exhibits endogeneity and environmental generalization capability over existing CDT. The simulation results prove the promising applications of PEM in 6G wireless networks.
II Implementation of Path Evolution Model
In Sec. II-A, we first design the intelligent operation framework of PEM. Next, environmental generalization advantages of PEM are analyzed in Sec. II-B. Then, a small amount of training data in practical situations can be achieved by efficient PEM learning approaches in Sec. II-C.
II-A PEM Framework
Figure 2: PEM operation framework, where the path extraction, path update and path evolution steps are illustrated in the bottom part.
The objective of the PEM operation framework is to construct, maintain and apply digital replicas of path evolutions upon existing communication systems, which is illustrated in Fig. 2. In existing communication systems, consecutive channel measurements can be attained by periodical pilot signals, e.g., Sounding Reference Signal (SRS). Then, the transmitted signals propagate from the UE to BS through multiple paths. Thus, the channel measurements contain the information of each path. Due to the universal spatial consistencies [7], path evolution exhibits obvious temporal correlation during UE mobility. Such consistencies enable continuous-time path evolution with discrete-time channel measurements. The procedure of PEM is included in the gray block of Fig. 2, which contains three main steps:
  1. 1.
    Path extraction: The objective of path extraction is to attain the feature of each path from input channel measurements, as illustrated in the bottom left of Fig. 2. For massive MIMO channel, each path can be characterized by the parameters of delay, AoA and power. Since the channel measurements are the sum of different paths, we first need to identify different paths with delay-angular representation or parameter estimation algorithms, e.g. SAGE and ESPRIT algorithms. In sparse-scattering environment, path parameters can be adopted as path features. In rich-scattering environment, multiple sub-paths with similar parameters form a path, which can be obtained via clustering. Then, path features can be represented as the power distribution in delay-angular domain.
  2. 2.
    Path update: Path update aims to refresh the temporal path dependencies with the latest extracted path features, which is the key to enabling the self-sustainability of PEM. As shown in the bottom middle of Fig. 2, it should first recognize the death of historical paths and the birth of the latest paths against channel non-stationarity. Next, the remaining latest path features can be associated with the surviving historical paths based on path feature discrepancies. Considering non-linearity in path feature evolutions, neural network is adopted to represent path feature temporal dependencies. The associated path features at current instant are fed into the neural network to update hidden states of each path, which refreshes path feature temporal dependencies.
  3. 3.
    Path evolution: The goal of path evolution is to attain the feature of each path at any desired instant via the updated temporal path feature dependencies, which is vital for real-time applications. The rationale lies in the fact that path features continuously vary between two adjacent update instants. Hereby, the non-linear path evolution is achieved by a well-trained neural network. With the previous path extraction step, the neural network for path evolution exhibits robust generalization capabilities among dynamic environments. Analysis of environmental generalization advantages is detailed in Sec. II-B.
Our proposed PEM operation framework is light-weighted and can be realized at a low cost. Firstly, PEM can avoid additional hardware devices and achieve protocol compatibility with existing communication systems. Secondly, in usual dynamic scenarios, the update period in PEM processing can be set in the scale of 0.1s, which is far longer than the typical slot length and achieves a low computation cost.
Figure 3: Illustration of PEM environmental generalizaition advantage analysis (top and middle), along with the efficient PEM learning approaches to reduce the amount of data collection under practical spatial and temporal constraints (bottom).
II-B Environmental Generalization
The invariance of target path evolutions to historical path features and time is fundamental for the PEM generalization capability. From block A to block E on the top of Fig. 3, analysis of invariant target path evolution of PEM is unfolded in a logically progressive manner as follows:
  1. 1.
    Block A: To begin with, we can focus on the generalizability in the feature evolution of a single path based on the path extraction step in the operation framework. Evolution of multiple paths in wireless channels includes two aspects, evolution of each single path and correlation among different paths. Through path extraction, the environment-related path correlation can be removed.
  2. 2.
    Block B: Subsequently, feature evolution of one single path originates from two physics processes, EM wave propagation and mobility. EM wave propagation (e.g., transmission and reflection) can be formulated as a function of interaction positions and EM parameters (e.g., conductivity and permittivity) alongside a path, which is universal among environments. The mobility model is a function of time and mobility parameters (e.g., initial interaction positions, velocity and acceleration), which affects the change of path feature. Without loss of generality, mobility model among environments can be unified [8].
  3. 3.
    Block C: Thereafter, original path evolution can be formulated as a function of time, EM and mobility parameters based on the EM wave propagation laws and mobility model.
  4. 4.
    Block D: Then, input in the original path evolution should be reformulated as historical path features to facilitate PEM. An auxiliary implicit expression of EM and mobility parameters can be attained from historical path features [9]. Since such implicit expression is determined by the EM propagation laws and mobility model, it is environment-generalizable as well.
  5. 5.
    Block E: Finally, due to the environment-generalizability in EM wave propagation laws and mobility model, the environmental generalizable target path evolution is yielded when the EM and mobility parameters are substituted by their implicit expressions.
Based on the invariance of target path evolutions, PEM exhibits better environmental generalization performance compared to existing CDT learning, as shown in the middle of Fig. 3. Under the framework of supervised learning, the trained neural network can approximate the underlying target function by minimizing training loss. Since the target path evolution of PEM is environment-generalizable, the trained neural network in PEM can also generalize to dynamic environments. On the contrary, target functions differ among environments in existing CDT learning. The rationale lies that the extrinsic input in existing CDTs does not include all the required information to directly obtain the desired output in dynamic environments. Thus, the target functions in existing CDTs are also controlled by environment-related factors (e.g., scatter layouts) of the training area. Consequently, the trained neural networks in existing CDTs degrade when they are directly deployed in dynamic environments.
II-C Efficient PEM Learning in Practical Situations
A small amount of training data in practical situations can be achieved with efficient PEM learning. For PEM learning, the trained neural network is yielded by minimizing loss function over a finite training dataset. Therefore, fitting accuracy of the trained neural network will be affected by the training sample distribution. Under practical constraints, the training dataset is collected within certain spatial and temporal ranges, which restricts its distribution. Fortunately, such constraints can be efficiently tackled with the following efficient PEM learning approaches.
Data augmentation is effective in achieving good fitting accuracy within a limited training area, which directly manipulates the training dataset and requires no additional training data in the new environments [10]. An example of data augmentation is illustrated in the bottom left of Fig. 3. Due to the limited spatial scale, the distribution range of the collected training samples is restricted. As a result, when neural networks trained from the original dataset are directly utilized, the performance of substantial out-of-distribution samples degrades. Fortunately, the distribution range of training samples can be easily broadened by data augmentation. As a result, when the neural network trained from the augmented training dataset is tested in a dynamic environment, its performance can still be guaranteed. Typical data augmentation techniques for path feature samples include scaling, rotation and translation.
In changing environments, continual learning strategies can reduce training data amount temporally [11]. An example of continual learning strategy is shown in the bottom right of Fig. 3. Since wireless environment gradually changes over time, sample distribution varies as well. To adapt to environment changes, we can continually collect training dataset and fine-tune neural networks accordingly. In addition to adapting to the latest environment, trained neural networks are also required to maintain the memories of old ones. To avoid the curse of forgetting, we can design loss functions when training in the latest environments, e.g., add network parameter regularization [11]. Since neural networks can reuse knowledge from previous environments, training samples amount can be reduced temporally with continual learning.
III Applications of Path Evolution Model
Once PEM is constructed and maintained, it can be applied in real time. As shown in the top of Fig. 4, two main applications of PEM can be categorized:
  • Multi-domain pilot overhead reduction: With temporal dependencies of path feature in PEM, frequency/time-domain ambiguity [12] for channel acquisition under low pilot densities can be tackled. Thus, pilot overhead in frequency/time-domain can be greatly reduced below channel coherence bandwidth and coherence time limitations. Additionally, beam prediction via path evolution can reduce pilot overhead in spatial-domain [13].
  • Multi-functional sensing: Note that explicit delay and angular information in path features characterize the geometrical relationship among BS, UE and scatters. Instant path features can be used for positioning and the shift of path features can be leveraged for velocity estimation. Meanwhile, with path trajectories maintained in PEM, user tracking can be achieved as well.
High-precision and low-overhead CSI acquisition capability of PEM is investigated by following numerical simulations.
Figure 4: Illustration of PEM applications (top), along with simulation scenario in Wireless Insite (bottom).
III-A Simulation Setup
In the simulations, precise ray-tracer Wireless Insite is adopted to generate CSI data. As shown in the bottom of Fig. 4, a street scenario is considered, where mobile users are distributed in two environments, namely, env-1 and env-2. Mobility of users is generated by microscopic SUMO simulator [14] with a 20 m/s speed limit, which leads to a maximum Doppler shift of 400 Hz. From the bottom of Fig. 4, it is apparent that environment properties distinctly differ between env-1 and env-2. For communication system configurations, bandwidth is 100 MHz and $16\times 1616 × 16$ Uniform Planar Arrays (UPA) are equipped in BS-1 and BS-2. Signal-to-Noise Ratio (SNR) of channel measurement is set as 10 dB. For the neural networks utilized in PEM and other deep learning based baselines, a training dataset with 1600 samples is collected from env-1. Data augmentation in Sec. II-C is adopted for efficient PEM learning.
Figure 5: An example of path evolution during UE mobility in env-1. Three physical paths are illustrated with a top view on the top side. Delay and AoA evolution of these paths are plotted on the bottom side: (a) path-1: LoS path; (b) path-2: reflection path from building-2; (c) path-3: reflection path from building-3.
Figure 6: Time-domain channel prediction under different SRS periods.
III-B PEM Functionality
Firstly, we evaluate the functionality of PEM by comparing the physical and digital path evolutions, which is illustrated in Fig. 5. Here, evolutions of three typical paths in env-1 are taken as examples, including one Line of Sight (LoS) path and two reflection paths. During UE mobility, it can be found that the digital evolutions of the three paths can accurately match their physical counterparts. Moreover, PEM can recognize the death of path-2 and the birth of path-3 when passing from building-2 to building-3. Thus, PEM can provide precise prior knowledge for CSI acquisition in real time. Additionally, the evolution error of path 3 in Fig. 5(c) can be progressively reduced with the updated channel measurements, which exhibits the self-sustainability of PEM.
III-C Time-domain Pilot Overhead Reduction
In this subsection, we aim to leverage PEM to reduce pilot overhead in time-domain by increasing SRS period. With high channel dynamics and large SRS period, continuous time-domain channel prediction is vital to guarantee seamless high spectral efficiency. Hereby, tensor neural Ordinary Differential Equation (ODE) channel prediction network [3] is adopted as a continuous CSI acquisition baseline without CDT. Normalized Mean Square Error (NMSE) of channel prediction schemes under different SRS periods is shown in Fig. 6. It can be found that NMSE of PEM is low and nearly remains unchanged when time-domain pilot overhead is reduced by 5 times. On the contrary, NMSE of the baseline obviously increases with the SRS period. Thus, it is evident that PEM can simultaneously enable low-overhead and high-precision CSI acquisition.
III-D Discussion over Generalization and Endogeneity
Intra-system endogeneity and environment generalization can be achieved in PEM. As depicted in Fig. 6, channel prediction performance of PEM is held in the unseen env-2, which is achieved without any external device. Such environmental generalization capability relies on the generalizability of target path evolution discussed in Sec. II-B and practical data augmentation techniques proposed in Sec. II-C. Additionally, we further consider a Channel Impulse Response (CIR) type CKM [15] as an existing CDT baseline, which achieves the NMSE of -15.37 dB in env-1 and 5.69 dB in env-2. Based on the simulation results, channel acquisition performances of these baselines are obviously degraded when tested in the unseen env-2 due to environment discrepancies, which shows the lack of generalizability.
IV Conclusions and Future Research
CDT is a new paradigm for efficient CSI acquisition in 6G wireless networks. In this paper, PEM is further introduced, which is designed to reflect path evolutions in the physical world. Different from existing vision information-based and CKM-based CDTs, PEM possesses three key advantages of full endogeneity, self-sustainability and environmental generalizability, which are vital to achieving low operation and deployment costs. To facilitate PEM, an intelligent operation framework is proposed, which is light-weighted and compatible with existing massive MIMO systems. The generalization advantages of PEM over other CDTs are rigorously analyzed and efficient PEM learning is proposed to reduce data collection burden practically. Based on extensive numerical simulations, we validate that high-precision and low-overhead CSI acquisition among dynamic environments can be effectively achieved by PEM, which can enable large system capacity for 6G wireless networks.
Although great potential and feasibility have been proved to adopt PEM for efficient CSI acquisition, its effectiveness relies on the specific dynamic nature of the wireless channel and the performance to process the input channel measurements during the operation. Hence, the future research directions span from environment-aware operation design to efficient data processing. Firstly, to optimize the operation performance of PEM, environment-aware update period should be designed by intelligent decision-making. Secondly, multi-band and multi-BS channel measurements are available in Frequency Division Duplexing (FDD) systems and dense networks, which enrich the information for PEM but also bring challenges due to their complicated internal correlations. Then, how to leverage the multi-band and multi-BS channel measurements deserves further investigation. Thirdly, PEM can tackle channel measurement noise with extrinsic information. However, the discrepancies among data modalities bring challenges to multi-modal data fusion. Thus, a multi-modal data fusion framework that fully employs specific physics meanings of extrinsic information requires further investigation.
V Acknowledgement
This work was supported by the National Key R&D Program of China under Grant 2022YFB2902004.

### v2

@@
 Abstract
-Massive Multiple Input Multiple Output (MIMO) is critical for boosting 6G wireless network capacity.
-Nevertheless, high dimensional Channel State Information (CSI) acquisition becomes the bottleneck of 6G massive MIMO system.
+Massive Multiple-Input Multiple-Output (MIMO) is critical for boosting 6G wireless network capacity.
+Nevertheless, high-dimensional Channel State Information (CSI) acquisition becomes the bottleneck of massive MIMO system.
 Recently, Channel Digital Twin (CDT), which replicates physical entities in wireless channels, has been proposed, providing site-specific prior knowledge for CSI acquisition.
-However, external devices (e.g., cameras and GPS devices) cannot always be integrated into existing communication systems, nor are they universally available across all scenarios.
+However, external devices cannot always be integrated into existing communication systems, nor are they universally available across all scenarios.
 Moreover, the trained CDT model cannot be directly applied in new environments, which lacks environmental generalizability.
 To this end, Path Evolution Model (PEM) is proposed as an alternative CDT to reflect physical path evolutions from consecutive channel measurements.
-Compared to existing CDTs, PEM demonstrates virtues of full endogeneity, self-sustainability and environmental generalizability.
+Compared to existing CDTs, PEM demonstrates virtues of full endogeneity, self-sustainability, and environmental generalizability.
 Firstly, PEM only requires existing channel measurements, which is free of other hardware devices and can be readily deployed.
 Secondly, self-sustaining maintenance of PEM can be achieved in dynamic channel by progressive updates.
-Thirdly, environmental generalizability can greatly reduce deployment costs in dynamic environments.
+Thirdly, environmental generalizability can greatly reduce costs for large-scale deployment.
 To facilitate the implementation of PEM, an intelligent and light-weighted operation framework is firstly designed.
-Then, the environmental generalizability of PEM is rigorously analyzed.
-Next, efficient learning approaches are proposed to reduce the amount of training data practically.
+Then, the environmental generalizability of PEM is rigorously analyzed, which can effectively address the distribution shift across environments.
 Extensive simulation results reveal that PEM can simultaneously achieve high-precision and low-overhead CSI acquisition, which can serve as a fundamental CDT for 6G wireless networks.
 I Introduction
-Massive Multiple Input Multiple Output (MIMO) is pivotal for 6G wireless networks, including ultra-high spectral efficiency, massive access and immersive sensing capabilities [1].
+Massive Multiple-Input Multiple-Output (MIMO) is pivotal for 6G wireless networks, including ultra-high spectral efficiency, massive access, and immersive sensing capabilities [11].
 To fully leverage large-scale antenna arrays, Channel State Information (CSI) knowledge is critical for beamforming and precoding operations in massive MIMO systems.
@@
 Consequently, effective throughput will not consistently grow with the number of antennas.
-High dimensional CSI acquisition becomes an inevitable challenge to meet promising 6G applications.
+High-dimensional CSI acquisition becomes an inevitable challenge to meet promising 6G applications.
 To support numerous 6G applications with high throughput, there are two main requirements for CSI acquisition: (1) high CSI precision to facilitate beamforming and precoding operations; (2) low acquisition overhead for simultaneously serving numerous User Equipments (UEs).
 I-A Channel Digital Twin
-To resolve the bottleneck of high dimensional CSI acquisition, mainstream approaches directly utilize the measured CSI for prediction, which can also be enhanced by deep learning [3].
-However, their channel acquisition accuracy dramatically relies on pilot densities, which cannot achieve high CSI precision when pilot densities are low.
-Additionally, deep learning related CSI extrapolation and interpolation heavily depend on the training dataset, whose generalization capability cannot be guaranteed in dynamic environments.
+To resolve the bottleneck of high-dimensional CSI acquisition, mainstream approaches directly utilize the measured CSI for prediction, which can be enhanced by deep learning [2].
+However, their channel acquisition accuracy dramatically relies on pilot densities, which cannot achieve high CSI precision with low pilot densities.
+Additionally, existing deep learning-related channel extrapolation and interpolation heavily depend on the distribution of the training dataset, which encounters the generalizability challenge when the distribution of wireless channel significantly shifts in the test environment.
 Recently, Channel Digital Twin (CDT) has emerged as an effective tool for CSI acquisition.
@@
 One is vision information, which provides geometric and object information to attain the properties of wireless channel [1, 14].
-Another is Channel Knowledge Map (CKM), which serves as a channel property database tagged with UE location [6].
-To infer wireless channel property (e.g., beam index and path parameters) with input data (e.g., images and UE locations), deep learning is usually adopted in current CDTs.
+Another is Channel Knowledge Map (CKM), which serves as a channel property database tagged with location [15].
+To infer wireless channel property (e.g., beam index and path parameters) with input data, deep learning is usually adopted in current CDTs.
 Despite the prior knowledge, current CDTs still encounter two challenges.
@@
 In summary, CDT is an effective tool to enable superior applications for 6G wireless networks.
-However, when considering the aforementioned two challenges, operation of current CDTs is still challenging.
+However, when considering the aforementioned two challenges, the operation of current CDTs is still challenging.
 Thus, the key question is can we find a new type of CDT that is endogenous within communication systems and is environment-generalizable?
@@
 As shown in the top left of Fig.
-1, the wireless channel between Base Station (BS) and UE is composited by multiple paths, which are basic physical entities for transmission, reflection and scattering.
-Due to UE mobility, feature (e.g., delay, Angle of Arrival (AoA), power) of each physical path consistently evolves with time, as shown in the top right of Fig.
+1, the wireless channel between the Base Station (BS) and UE is composited by multiple paths, which are basic physical entities under transmission, reflection, and scattering.
+Due to UE mobility, the path feature (e.g., path parameters and power distribution in the delay-angular domain) of each physical path consistently evolves with time, as shown in the top right of Fig.
 1.
 Since channel measurements are frequently conducted in existing communication systems, as shown at the bottom of Fig.
-1, the concept of PEM aims to extract each physical path and then attain digital replicas of their feature temporal evolutions from the consecutive channel measurements.
+1, the concept of PEM aims to extract the paths and then attain digital replicas of their feature temporal evolutions from the consecutive channel measurements.
 Then, CDT is yielded when digital replicas of all path evolutions are summed up.
-Thus, the constructed PEM can be applied to directly attain the desired path feature at any instant, which greatly reduces pilot overhead and enables the functionality of CDT.
+Thus, the constructed PEM can be applied to directly attain the desired path feature at future instants, which greatly reduces pilot overhead and enables the functionality of CDT.
 Firstly, the processing of PEM only requires existing channel measurements from communication systems, which is fully endogenous.
 Secondly, path feature evolutions in PEM are determined by universal electromagnetic (EM) propagation principles and regular user mobility.
-As a result, a generalizable path feature evolution pattern can be assumed among dynamic environments.
-Therefore, when a neural network is trained to fit target path feature evolution, it can generalize to different environments as well.
+As a result, a generalizable path feature evolution can be assumed among dynamic environments, which facilitates environment-generalizability.
 Figure 1: Structure of PEM, where digital replicas of physical path evolutions are created.
 I-C Advantages
-Proposed PEM exhibits three inherent advantages over other existing CDTs:
+The proposed PEM exhibits three inherent advantages over other existing CDTs:
   1. 1.
-    Fully-endogenous: The intra-system endogeneity of PEM has two folds of virtues.
-On the one hand, only existing channel measurements within communication systems are required, which is free of other external devices.
-On the other hand, communication signals can realize larger coverage than either light or GPS signals.
-As a result, the outage probability of PEM can be much lower than the existing vision information-based and CKM-based CDTs.
+    Fully-endogenous: PEM can be operated without introducing external hardware devices, which can be seamlessly integrated into existing communication systems.
   2. 2.
@@
   3. 3.
-    Environment-generalizable: Trained neural network in PEM shows remarkable environmental generalizability since it endeavors to fit the environment-generalizable target path feature evolution.
+    Environment-generalizable: With efficient learning, the neural network in PEM exhibits strong environmental generalizability.
 Such environmental generalizability greatly relieves the data collection and model retraining burden in new environments, which enables large-scale PEM deployment in 6G wireless networks.
@@
 In this paper, we propose the concept of PEM, aiming to create digital replicas of physical path evolutions upon communication systems.
-Firstly, we propose an intelligent operation framework to efficiently construct, maintain and apply PEM in dynamic wireless channels, which is compatible with existing communication systems and can be operated at low cost.
-Secondly, the environmental generalization capability of PEM is carefully analyzed and compared with existing CDT learning.
-Thirdly, efficient learning approaches are proposed to maximize the robustness of PEM under practical data collection constraints.
-In the numerical simulations, a good match can be found by comparing the physical and digital path evolutions, which validate the functionality of the proposed PEM.
+Firstly, we propose an intelligent operation framework to efficiently construct, maintain, and apply PEM in dynamic wireless channels, which is compatible with existing communication systems and can be operated at a low cost.
+Secondly, the environment invariance of the target function of PEM is presented, which is fundamental to achieving environment generalizability.
+Thirdly, distribution shifts across environments are progressively addressed to enable environment generalizability for PEM.
+In the simulations, a good match can be found by comparing the physical and digital path evolutions, which validate the functionality of the proposed PEM.
 To justify the pilot overhead reduction capability of PEM, channel prediction under different time-domain pilot densities is investigated.
-Compared to CSI acquisitions without CDT, both high-precision and low-overhead CSI acquisition can be achieved with PEM.
-In addition, PEM exhibits endogeneity and environmental generalization capability over existing CDT.
-The simulation results prove the promising applications of PEM in 6G wireless networks.
+Compared to CSI acquisition without CDT, both high-precision and low-overhead CSI acquisition can be achieved with PEM.
+Additionally, PEM exhibits endogeneity and strong environmental generalization capability.
 II Implementation of Path Evolution Model
-In Sec.
-II-A, we first design the intelligent operation framework of PEM.
-Next, environmental generalization advantages of PEM are analyzed in Sec.
-II-B.
-Then, a small amount of training data in practical situations can be achieved by efficient PEM learning approaches in Sec.
-II-C.
 II-A PEM Framework
-Figure 2: PEM operation framework, where the path extraction, path update and path evolution steps are illustrated in the bottom part.
-The objective of the PEM operation framework is to construct, maintain and apply digital replicas of path evolutions upon existing communication systems, which is illustrated in Fig.
+The objective of the PEM operation framework is to construct, maintain, and apply digital replicas of path evolutions upon existing communication systems, which is illustrated in Fig.
 2.
 In existing communication systems, consecutive channel measurements can be attained by periodical pilot signals, e.g., Sounding Reference Signal (SRS).
-Then, the transmitted signals propagate from the UE to BS through multiple paths.
+Then, the transmitted signals propagate from the UE to the BS through multiple paths.
 Thus, the channel measurements contain the information of each path.
@@
 Such consistencies enable continuous-time path evolution with discrete-time channel measurements.
-The procedure of PEM is included in the gray block of Fig.
+The procedure of PEM is demonstrated in the gray block of Fig.
 2, which contains three main steps:
+Figure 2: PEM operation framework, where the path extraction, path update and path evolution steps are illustrated in the bottom part.
   1. 1.
@@
 2.
-For massive MIMO channel, each path can be characterized by the parameters of delay, AoA and power.
+For massive MIMO channel, each path can be characterized by the parameters of delay, Angle of Arrival (AoA) and power.
 Since the channel measurements are the sum of different paths, we first need to identify different paths with delay-angular representation or parameter estimation algorithms, e.g.
 SAGE and ESPRIT algorithms.
-In sparse-scattering environment, path parameters can be adopted as path features.
-In rich-scattering environment, multiple sub-paths with similar parameters form a path, which can be obtained via clustering.
-Then, path features can be represented as the power distribution in delay-angular domain.
+In the sparse-scattering environment, path parameters can be adopted as path features.
+In the rich-scattering environment, multiple sub-paths with similar parameters form a path, which can be obtained via clustering.
+Then, path features can be represented as the power distribution in the delay-angular domain.
+The accuracy of path extraction is determined by the measurement resources, including bandwidth and the number of antennas.
   2. 2.
@@
 As shown in the bottom middle of Fig.
-2, it should first recognize the death of historical paths and the birth of the latest paths against channel non-stationarity.
-Next, the remaining latest path features can be associated with the surviving historical paths based on path feature discrepancies.
+2, PEM can first recognize the birth/death status of paths by comparing the path feature distance with a pre-defined threshold.
+Thus, the proposed PEM is robust to the blockage scenarios by terminating the evolution of the disappearing Line of Sight (LoS) path.
+Next, the remaining latest path features can be associated with the surviving historical paths based on path feature distances, which can be solved by bipartite matching.
 Considering non-linearity in path feature evolutions, neural network is adopted to represent path feature temporal dependencies.
-The associated path features at current instant are fed into the neural network to update hidden states of each path, which refreshes path feature temporal dependencies.
+The associated path features at the current instant are fed into the neural network to update the hidden states of each path, which refreshes path feature temporal dependencies.
   3. 3.
@@
 Hereby, the non-linear path evolution is achieved by a well-trained neural network.
-With the previous path extraction step, the neural network for path evolution exhibits robust generalization capabilities among dynamic environments.
-Analysis of environmental generalization advantages is detailed in Sec.
-II-B.
+With the previous path extraction step, the neural network for path evolution exhibits robust generalization capabilities across environments.
+Analysis of environmental generalizability is detailed in Sec.
+II-B and II-C.
 Our proposed PEM operation framework is light-weighted and can be realized at a low cost.
 Firstly, PEM can avoid additional hardware devices and achieve protocol compatibility with existing communication systems.
-Secondly, in usual dynamic scenarios, the update period in PEM processing can be set in the scale of 0.1s, which is far longer than the typical slot length and achieves a low computation cost.
-Figure 3: Illustration of PEM environmental generalizaition advantage analysis (top and middle), along with the efficient PEM learning approaches to reduce the amount of data collection under practical spatial and temporal constraints (bottom).
-II-B Environmental Generalization
+Secondly, considering the average lifetime of paths in usual dynamic scenarios, the update period in PEM processing can be set in the scale of 0.1s, which is far longer than the typical slot length and makes the pilot overhead affordable.
+Thirdly, the historical path feature sequence can be encoded in the hidden states of the neural network of PEM.
+Thus, PEM only needs to retain the extracted path feature at the current instant, which makes the memory occupation affordable.
+Figure 3: Analysis of environment-invariant target function in PEM is shown on the top.
+Then, an overview environment generalizable PEM is illustrated in the middle, which progressively addresses the distribution shift.
+Next, data augmentation and continual learning that address the varying user mobility are detailed at the bottom.
+II-B Environment Invariance of Target Function
 The invariance of target path evolutions to historical path features and time is fundamental for the PEM generalization capability.
 From block A to block E on the top of Fig.
-3, analysis of invariant target path evolution of PEM is unfolded in a logically progressive manner as follows:
+3, analysis of invariant target path evolution of PEM is unfolded in a logically progressive manner below.
   1. 1.
     Block A: To begin with, we can focus on the generalizability in the feature evolution of a single path based on the path extraction step in the operation framework.
-Evolution of multiple paths in wireless channels includes two aspects, evolution of each single path and correlation among different paths.
+The evolution of multiple paths in wireless channels includes two aspects: the evolution of each single path and the correlation among different paths.
 Through path extraction, the environment-related path correlation can be removed.
   2. 2.
-    Block B: Subsequently, feature evolution of one single path originates from two physics processes, EM wave propagation and mobility.
-EM wave propagation (e.g., transmission and reflection) can be formulated as a function of interaction positions and EM parameters (e.g., conductivity and permittivity) alongside a path, which is universal among environments.
-The mobility model is a function of time and mobility parameters (e.g., initial interaction positions, velocity and acceleration), which affects the change of path feature.
+    Block B: Subsequently, the evolution of a single path originates from two physics processes, EM wave propagation and mobility.
+EM wave propagation (e.g., transmission and reflection) can be formulated as a function of interaction positions and EM parameters (e.g., conductivity and permittivity) along a path, which is universal among environments.
+The user mobility model is calculated based on kinematics, which can be parameterized by initial conditions, speed limit, and topology nodes.
 Without loss of generality, mobility model among environments can be unified [3].
   3. 3.
-    Block C: Thereafter, original path evolution can be formulated as a function of time, EM and mobility parameters based on the EM wave propagation laws and mobility model.
+    Block C: Thereafter, original path evolution can be formulated as a function of time, EM, and mobility parameters based on the EM wave propagation laws and mobility model.
   4. 4.
     Block D: Then, input in the original path evolution should be reformulated as historical path features to facilitate PEM.
-An auxiliary implicit expression of EM and mobility parameters can be attained from historical path features [9].
-Since such implicit expression is determined by the EM propagation laws and mobility model, it is environment-generalizable as well.
+With sufficient measurements and Signal-to-Noise Ratio (SNR), an auxiliary implicit expression of EM and mobility parameters can be attained from historical path features [4].
+Since the implicit expression is determined by the EM propagation laws and mobility model, it is environment-invariant as well.
   5. 5.
-    Block E: Finally, due to the environment-generalizability in EM wave propagation laws and mobility model, the environmental generalizable target path evolution is yielded when the EM and mobility parameters are substituted by their implicit expressions.
-Based on the invariance of target path evolutions, PEM exhibits better environmental generalization performance compared to existing CDT learning, as shown in the middle of Fig.
+    Block E: Finally, due to the invariance of EM wave propagation laws and mobility model, the environment-invariant target path evolution is yielded when the EM and mobility parameters are substituted by their implicit expressions.
+II-C Addressing Distribution Shift Across Environments
+Based on the invariance of the target function, distribution shifts across environments should be addressed to enable environment generalizability for PEM.
+The distribution shift of wireless channel evolution during user mobility is composed of three factors: (1) the distribution shift of multipath structure; (2) the distribution shift of instant path feature; (3) the distribution shift of path feature variation.
+Firstly, the multipath structure includes the number of paths and the dependency among paths.
+Due to the diversity of shape, deployment, and EM properties of objects in wireless channels, the distribution of multipath structure significantly shifts across different environments.
+Secondly, the distribution of the instant path feature is determined by the interaction points along a path.
+Due to the variations of object configurations and user positions, the distribution of instant path feature diverges in different environments as well.
+Thirdly, path feature variation is caused by user mobility, where distribution shift occurs when user mobility pattern changes.
+The proposed PEM can progressively address the aforementioned three types of distribution shift, which is shown in the middle of Fig.
 3.
-Under the framework of supervised learning, the trained neural network can approximate the underlying target function by minimizing training loss.
-Since the target path evolution of PEM is environment-generalizable, the trained neural network in PEM can also generalize to dynamic environments.
-On the contrary, target functions differ among environments in existing CDT learning.
-The rationale lies that the extrinsic input in existing CDTs does not include all the required information to directly obtain the desired output in dynamic environments.
-Thus, the target functions in existing CDTs are also controlled by environment-related factors (e.g., scatter layouts) of the training area.
-Consequently, the trained neural networks in existing CDTs degrade when they are directly deployed in dynamic environments.
-II-C Efficient PEM Learning in Practical Situations
-A small amount of training data in practical situations can be achieved with efficient PEM learning.
-For PEM learning, the trained neural network is yielded by minimizing loss function over a finite training dataset.
-Therefore, fitting accuracy of the trained neural network will be affected by the training sample distribution.
-Under practical constraints, the training dataset is collected within certain spatial and temporal ranges, which restricts its distribution.
-Fortunately, such constraints can be efficiently tackled with the following efficient PEM learning approaches.
-Data augmentation is effective in achieving good fitting accuracy within a limited training area, which directly manipulates the training dataset and requires no additional training data in the new environments [10].
-An example of data augmentation is illustrated in the bottom left of Fig.
-3.
-Due to the limited spatial scale, the distribution range of the collected training samples is restricted.
-As a result, when neural networks trained from the original dataset are directly utilized, the performance of substantial out-of-distribution samples degrades.
-Fortunately, the distribution range of training samples can be easily broadened by data augmentation.
-As a result, when the neural network trained from the augmented training dataset is tested in a dynamic environment, its performance can still be guaranteed.
-Typical data augmentation techniques for path feature samples include scaling, rotation and translation.
-In changing environments, continual learning strategies can reduce training data amount temporally [11].
-An example of continual learning strategy is shown in the bottom right of Fig.
-3.
-Since wireless environment gradually changes over time, sample distribution varies as well.
-To adapt to environment changes, we can continually collect training dataset and fine-tune neural networks accordingly.
-In addition to adapting to the latest environment, trained neural networks are also required to maintain the memories of old ones.
-To avoid the curse of forgetting, we can design loss functions when training in the latest environments, e.g., add network parameter regularization [11].
-Since neural networks can reuse knowledge from previous environments, training samples amount can be reduced temporally with continual learning.
+Firstly, the neural network in PEM is designed in a path-oriented manner.
+Explicitly, the neural network individually realizes the evolution of each extracted path, which tackles the distribution shift of the multipath structure.
+Secondly, path feature variation learning is adopted.
+Hereby, historical path feature variations are calculated and input into the neural network, which predicts path feature variation in future instants.
+Then, the predicted path feature can be obtained by adding the predicted path feature variation and the latest historical path feature.
+For the LoS path and specular reflection paths, the variation of path features weakly correlates with the instant path feature, which largely mitigates the distribution shift of the instant path feature.
+Thirdly, data augmentation and continual learning are adopted to address the distribution shift of user mobility pattern.
+Compared to the object deployment and materials changes, user mobility exhibits a much lower variability across environments, which facilitates environment generalization.
+As shown in the bottom left of Fig.
+3, data augmentation directly manipulates the training samples, which can enhance the diversity of the training dataset and is free of training samples from the unseen environment [12].
+Typical data augmentation includes rotation and scaling, which benefits the generalization to new environments with different speed ranges and orientations.
+As depicted in the bottom right of Fig.
+3, continual learning can be leveraged to adapt to the variation of user mobility pattern, which will be triggered when a large amount of out-of-distribution mobile users occur.
+Note that the in-distribution interval of user mobility pattern is far greater than the operation duration of a single user, data collection and model update in continual learning are tractable.
+Meanwhile, since the memories of seen environments can be retained under a continual learning framework [13], the generalizability to new environments in PEM can also be gradually increased.
 III Applications of Path Evolution Model
@@
   • Multi-domain pilot overhead reduction: With temporal dependencies of path feature in PEM, frequency/time-domain ambiguity [10] for channel acquisition under low pilot densities can be tackled.
-Thus, pilot overhead in frequency/time-domain can be greatly reduced below channel coherence bandwidth and coherence time limitations.
+Thus, pilot overhead in the frequency/time-domain can be greatly reduced below channel coherence bandwidth and coherence time limitations.
 Additionally, beam prediction via path evolution can reduce pilot overhead in spatial-domain [8].
-  • Multi-functional sensing: Note that explicit delay and angular information in path features characterize the geometrical relationship among BS, UE and scatters.
+  • Multi-functional sensing: Note that explicit delay and angular information in path features characterize the geometrical relationship among BS, UE, and scatters.
 Instant path features can be used for positioning and the shift of path features can be leveraged for velocity estimation.
 Meanwhile, with path trajectories maintained in PEM, user tracking can be achieved as well.
-High-precision and low-overhead CSI acquisition capability of PEM is investigated by following numerical simulations.
+The high-precision and low-overhead CSI acquisition capability of PEM is investigated by the following simulations.
 Figure 4: Illustration of PEM applications (top), along with simulation scenario in Wireless Insite (bottom).
 III-A Simulation Setup
-In the simulations, precise ray-tracer Wireless Insite is adopted to generate CSI data.
+In the simulations, the precise ray-tracer Wireless Insite is adopted to generate CSI data.
 As shown in the bottom of Fig.
-4, a street scenario is considered, where mobile users are distributed in two environments, namely, env-1 and env-2.
-Mobility of users is generated by microscopic SUMO simulator [14] with a 20 m/s speed limit, which leads to a maximum Doppler shift of 400 Hz.
-From the bottom of Fig.
-4, it is apparent that environment properties distinctly differ between env-1 and env-2.
-For communication system configurations, bandwidth is 100 MHz and $16\times 1616 × 16$ Uniform Planar Arrays (UPA) are equipped in BS-1 and BS-2.
-Signal-to-Noise Ratio (SNR) of channel measurement is set as 10 dB.
-For the neural networks utilized in PEM and other deep learning based baselines, a training dataset with 1600 samples is collected from env-1.
+4, street scenarios are considered, where mobile users are distributed in env-1 and env-2.
+Explicitly, env-1 and env-2 exhibit distinct geometric layouts, including building layouts, BS heights, and user positions relative to BS.
+Additionally, different types of materials are adopted in the buildings in env-1 and env-2.
+Mobility of users is generated by SUMO simulator [6], where speed limits are set as 20 m/s and 30 m/s in env-1 and env-2, respectively.
+The communication systems operate at a 6 GHz carrier frequency, where the bandwidth is set as 100 MHz and the number of subcarriers is 138.
+Uniform Planar Arrays (UPA) with shape $16\times 16$ are both equipped in BS-1 and BS-2.
+The SNR of channel measurement is set as 10 dB.
+For the neural network in PEM and deep learning-based baseline, a training dataset with 1600 samples is collected from env-1, which is free of training samples in env-2.
 Data augmentation in Sec.
-II-C is adopted for efficient PEM learning.
+II-C is adopted during the training of PEM.
 Figure 5: An example of path evolution during UE mobility in env-1.
@@
 Delay and AoA evolution of these paths are plotted on the bottom side: (a) path-1: LoS path; (b) path-2: reflection path from building-2; (c) path-3: reflection path from building-3.
-Figure 6: Time-domain channel prediction under different SRS periods.
 III-B PEM Functionality
@@
 5.
-Here, evolutions of three typical paths in env-1 are taken as examples, including one Line of Sight (LoS) path and two reflection paths.
+Here, evolutions of three typical paths in env-1 are taken as examples, including one LoS path and two reflection paths.
 During UE mobility, it can be found that the digital evolutions of the three paths can accurately match their physical counterparts.
@@
 III-C Time-domain Pilot Overhead Reduction
-In this subsection, we aim to leverage PEM to reduce pilot overhead in time-domain by increasing SRS period.
-With high channel dynamics and large SRS period, continuous time-domain channel prediction is vital to guarantee seamless high spectral efficiency.
-Hereby, tensor neural Ordinary Differential Equation (ODE) channel prediction network [3] is adopted as a continuous CSI acquisition baseline without CDT.
+In this subsection, we aim to leverage PEM to reduce pilot overhead in time-domain by increasing the SRS period.
+With high channel dynamics and a large SRS period, continuous time-domain channel prediction is vital to guarantee seamless high spectral efficiency.
+Hereby, tensor neural Ordinary Differential Equation (ODE) channel prediction network [2] and model-based prediction [9] is adopted as a continuous CSI prediction baseline without CDT, where the length of the historical CSI sequence is 8.
 Normalized Mean Square Error (NMSE) of channel prediction schemes under different SRS periods is shown in Fig.
 6.
-It can be found that NMSE of PEM is low and nearly remains unchanged when time-domain pilot overhead is reduced by 5 times.
-On the contrary, NMSE of the baseline obviously increases with the SRS period.
+It can be found that the NMSE of PEM is low and nearly remains unchanged when time-domain pilot overhead is reduced by 5 times, which benefits from the real-time prior knowledge of path feature evolution.
+Caused by the time-domain ambiguity, the NMSE of the baselines obviously increases when the SRS period is large.
 Thus, it is evident that PEM can simultaneously enable low-overhead and high-precision CSI acquisition.
@@
 As depicted in Fig.
-6, channel prediction performance of PEM is held in the unseen env-2, which is achieved without any external device.
-Such environmental generalization capability relies on the generalizability of target path evolution discussed in Sec.
-II-B and practical data augmentation techniques proposed in Sec.
-II-C.
-Additionally, we further consider a Channel Impulse Response (CIR) type CKM [15] as an existing CDT baseline, which achieves the NMSE of -15.37 dB in env-1 and 5.69 dB in env-2.
-Based on the simulation results, channel acquisition performances of these baselines are obviously degraded when tested in the unseen env-2 due to environment discrepancies, which shows the lack of generalizability.
+6, the channel prediction performance of PEM is held in the unseen env-2, which is achieved without any external device.
+Note that 20% of the user distribution area is blocked in env-2.
+Thus, the simulation results reveal that the PEM is robust to blockage.
+Conversely, the channel acquisition performance of the deep learning-based baseline is obviously degraded when tested in the unseen env-2, which shows the lack of generalizability.
+Figure 6: Time-domain channel prediction under different SRS periods.
 IV Conclusions and Future Research
 CDT is a new paradigm for efficient CSI acquisition in 6G wireless networks.
-In this paper, PEM is further introduced, which is designed to reflect path evolutions in the physical world.
-Different from existing vision information-based and CKM-based CDTs, PEM possesses three key advantages of full endogeneity, self-sustainability and environmental generalizability, which are vital to achieving low operation and deployment costs.
+In this paper, PEM is introduced as a new type of CDT, which is designed to reflect path evolutions in the physical world.
+Compared to existing vision information-based and CKM-based CDTs, PEM possesses three key advantages of full endogeneity, self-sustainability, and environmental generalizability, which are vital to achieving low operation and deployment costs.
 To facilitate PEM, an intelligent operation framework is proposed, which is light-weighted and compatible with existing massive MIMO systems.
-The generalization advantages of PEM over other CDTs are rigorously analyzed and efficient PEM learning is proposed to reduce data collection burden practically.
-Based on extensive numerical simulations, we validate that high-precision and low-overhead CSI acquisition among dynamic environments can be effectively achieved by PEM, which can enable large system capacity for 6G wireless networks.
+The environment generalizability of PEM is rigorously analyzed, which is realized by progressively addressing the distribution shift.
+Based on extensive simulations, we validate that high-precision and low-overhead CSI acquisition among dynamic environments can be effectively achieved by PEM, which can enable large system capacity for 6G wireless networks.
 Although great potential and feasibility have been proved to adopt PEM for efficient CSI acquisition, its effectiveness relies on the specific dynamic nature of the wireless channel and the performance to process the input channel measurements during the operation.
 Hence, the future research directions span from environment-aware operation design to efficient data processing.
-Firstly, to optimize the operation performance of PEM, environment-aware update period should be designed by intelligent decision-making.
-Secondly, multi-band and multi-BS channel measurements are available in Frequency Division Duplexing (FDD) systems and dense networks, which enrich the information for PEM but also bring challenges due to their complicated internal correlations.
-Then, how to leverage the multi-band and multi-BS channel measurements deserves further investigation.
+Firstly, to optimize the operation performance of PEM, the environment-aware update period should be designed by intelligent decision-making.
+Secondly, the accuracy of path extraction in PEM is affected by the measurement resources.
+With limited measurement resources, how to achieve accurate path extraction in a rich multipath environment needs to be resolved.
 Thirdly, PEM can tackle channel measurement noise with extrinsic information.
@@
 Thus, a multi-modal data fusion framework that fully employs specific physics meanings of extrinsic information requires further investigation.
+Fourthly, in complex environments with diffusion paths and moving scatters, the distribution shift of the path features evolution becomes more complicated.
+Therefore, efficient processing algorithms in PEM based on a more comprehensive distribution shift model merit a deeper exploration.
 V Acknowledgement